OpenStudy (anonymous):

what is the slope of a line perpendicular to the line 6x + 3y = 15

jigglypuff314 (jigglypuff314):

the slope of a line perpendicular to a line would be the flipped opposite of the given slope. so for Example, if you were given a slope of 4/3 the slope of the perpendicular would be -3/4

OpenStudy (anonymous):

how would i show work

geerky42 (geerky42):

just show your work of turning given equation into slope-intercept form equation.

geerky42 (geerky42):

then take slope (m) and turn it into -1/m

geerky42 (geerky42):

if slope is m, then the slope of line perpendicular of original line is -1/m, just like example @jigglypuff314 gave you.

OpenStudy (anonymous):

i dont understand what u guys are saying :'( . i did an additive inverse

jigglypuff314 (jigglypuff314):

ok, well first figure out the slope of the given equation :) do you know how to do that?

OpenStudy (anonymous):

i know like the formula for slope: y2-y1/x2-x1

OpenStudy (anonymous):

right?

jigglypuff314 (jigglypuff314):

that's when you're given two points :) for when you're given an equation, you would want to get it into slope intercept form y = mx + b where m = slope so solve for y from 6x + 3y = 15

OpenStudy (anonymous):

oh. so

OpenStudy (anonymous):

3y= 6x + 15?

jigglypuff314 (jigglypuff314):

3y = -6x + 15 :) good, then divide both sides by 3 to get y alone

OpenStudy (anonymous):

so y = -2x + 5

jigglypuff314 (jigglypuff314):

yep! :D then given y = mx + b where m=slope find the slope of y = -2x + 5

OpenStudy (anonymous):

-2? :D

jigglypuff314 (jigglypuff314):

correct! :) now take the flipped opposite of it like I showed in the first comment

OpenStudy (anonymous):

umm. idk what that means ._.

jigglypuff314 (jigglypuff314):

for Example, if you were given a slope of 4/3 the slope of the perpendicular would be -3/4

OpenStudy (anonymous):

so. the slope is positive 2?

jigglypuff314 (jigglypuff314):

close ^_^ don't forget to flip it

OpenStudy (anonymous):

1/2 ?

jigglypuff314 (jigglypuff314):

Correct!!! :D :D :D

OpenStudy (anonymous):

YAY